OpenStudy (anonymous):

real square roots of 225

OpenStudy (anonymous):

\[\pm 15\]

OpenStudy (anonymous):

what does this mean real square roots? square roots are positive. that is what it means "square root" of something. it is not the same as saying solve for x \[x^2=225\] where the answer is \[\pm 15\]

OpenStudy (anonymous):

if they are teaching you that there are two square roots of 225 they are mistaken. there are two solutions to \[x^2=225\] but there is only one \[\sqrt{225}\]

OpenStudy (anonymous):

^ Depends on who you ask :D When they say "real," I go for ALL real answers :D